OpenStudy (anonymous):

Write the sum using summation notation, assuming the suggested pattern continues. -10 - 2 + 6 + 14 + ... + 110

OpenStudy (anonymous):

@iambatman

OpenStudy (anonymous):

@zepdrix

OpenStudy (anonymous):

@ParthKohli

OpenStudy (jango_in_dtown):

Hii @lxoser

OpenStudy (jango_in_dtown):

see here first term =-10 second term =-2=-10+8 third term =6=-10+2*8

OpenStudy (jango_in_dtown):

let the nth term be 110 see the first term is -10 and common difference is 8 then -10+(n-1)8=110 implies (n-1)=15 implies n=16 so the required answer is

OpenStudy (jango_in_dtown):

\[\sum_{n=1}^{16}(-10+(n-1)*8)\]

OpenStudy (anonymous):

these are my options.. the one highlighted looks close to your answer.

OpenStudy (jango_in_dtown):

@lxoser both are equivalent. if you put n in case of n-1, then the lower limit will change to 0/// yes it is correct